OpenStudy (anonymous):

x^2-16/x-9 . x^2+x-90/x^2+14x+40

OpenStudy (jdoe0001):

\(\large \begin{array}{cccl} x^2&+x&-90\\ &\uparrow &\uparrow \\ &10-9&10\cdot -9 \end{array}\qquad \begin{array}{cccll} x^2&+14x&+40\\ &\uparrow &\uparrow \\ &10+4&10\cdot 4 \end{array}\) can you factor those 2 fellows out? what do you think?